OpenStudy (anonymous):

-3+a=-5 solve using the addition principle

OpenStudy (amorfide):

add 3 on both sides to get a=-5+3 a=-2

OpenStudy (anonymous):

-3+a=-5 -3 + 3 + a = -5 + 3 a = -2

OpenStudy (wolfboy):

I agree with both of them and how they did it